OpenStudy (anonymous):

what is the center of the following circle? (x+12)^2+(y+6)^2=64

OpenStudy (anonymous):

center is (-12,-6)

OpenStudy (anonymous):

Help plz :C Thankew

OpenStudy (kohai):

The center of a circle is (h, k). Your current equation is basically this: (x – h)^2 + (y – k)^2 = r2 So your center will be (-12, -6).

OpenStudy (jdoe0001):

@UserA113 http://www.math-aids.com/images/pythagorean-definition-05.png <-- use the pythagorean theorem

OpenStudy (jdoe0001):

\(\bf (x-{\color{blue}{ h}})^2+(y-{\color{brown}{ k}})^2=r^2\qquad center\ ({\color{blue}{ h}},{\color{brown}{ k}})\)

OpenStudy (anonymous):

ikk i tried its just confusingg

OpenStudy (jdoe0001):

 \(\large \bf c^2=a^2+b^2\implies \sqrt{c^2-a^2}=b\)

OpenStudy (anonymous):

is that the formula

OpenStudy (jdoe0001):

for the pythagorean theorem. yes

OpenStudy (anonymous):

my question is dif tho its like 117 x+3 and x

OpenStudy (jdoe0001):

the picture doesn't look like that

OpenStudy (anonymous):

b=35.78=16 root 5 Which is C